1. Calculate the moles of glyceryl tripalmitate. Answer with 5 significant figures

2. Calculate the expected moles of sodium palmititate formed. Answer with 5 significant figures.

3. Calculate the theoretical yield of the soap. Answer with 5 significant figures.

4. Calculate the percent yield of the soap. Answer with 4 significant figures

Mass of Soap - 18.965g

Mass of NaOH- 10.122g

Mass of Oil- 10.230g
Answer #1

Q1:

Number of moles of glyceryl tripalmitate = 10.230g/807.339g/mol = 0.012671 mol

Q2:

Number of moles of palmitic acid = 0.012671mol × 3 = 0.038014 mol

Number of moles of NaOH = 10.122g/40g/mol = 0.25305mol

Expected moles of sodium palmitate formed = 0.038014 mol

Q3:

Number of moles of sodium tripalmitate formed = 0.038014mol

Theoretical yield of soap = 0.038014mol×278.411g/mol = 10.583g

Q4:

% yield = 18.965g/10.583g ×100 % = 179.2 %